OpenStudy (anonymous):

guys how to solve this (tan^3)x help me guys with solution :(

OpenStudy (perl):

theres no equation?

OpenStudy (anonymous):

oh imean integrate of (tan^3)x

OpenStudy (anonymous):

You know that tan²x=sec²x−1. Also d(tanx)/dx=sec²x, you you may get something. Lets try it out. ∫tan³x dx= ∫tan²x tanx dx =∫(sec²x−1)tanx dx =∫sec²xtanx dx − ∫tanx dx Put u=tanx, then du/dx=sec²x, du=sec²xdx =∫u du − ∫tanx dx =u²/2 − log|secx| + C =(tan²x)/2 − log|secx| + C

OpenStudy (anonymous):

from yahoo

OpenStudy (anonymous):

do you think is that a correct?

OpenStudy (anonymous):

oh i see.. thank you so much

OpenStudy (abb0t):

Wait, why are you doing integration? I don't see anything about integration being asked here. Lol